OpenStudy (anonymous):

Given h = -2 and k = 4, what is the equation of the graph if the parent function is y=√x? Describe the transformation.

OpenStudy (danjs):

\[y = \sqrt{(x-h)} + k\] \[y = \sqrt{x + 2} + 4\] that what you are looking for, translations?

OpenStudy (anonymous):

yes tahts perfect thank u

OpenStudy (danjs):

moves it left 2 units and up 4 units